a. Show that the curve with parametric equations

x=sint and y=sin(t+sint) for 0≤t≤2π

has two tangent lines at the origin and their equations. Illustrate by graphing the curve and its tangents.


b. Find the slope of the line tangent to the parametric curve

x=2cost and y=2-cos²t for 0≤t≤π

at points (-1, -1). Show the graph the parametric equations and the tangent line.


1
Expert's answer
2021-11-22T15:52:48-0500

a.


"\begin{matrix}
 x=0 \\
 y=0
\end{matrix}, \ 0\leq t\leq \pi"

"\begin{matrix}
 \sin t=0 \\
 \sin(t+\sin t)=0
\end{matrix}, \ 0\leq t\leq \pi"

"Origin(0, 0): t=0\ or\ t=\pi"

"slope=\dfrac{dy}{dx}=\dfrac{dy/dt}{dx/dt}=\dfrac{\cos(t+\sin t)\cdot(1+\cos t)}{\cos t}"

"t_1=0:"


"slope_1=\dfrac{\cos((0)+\sin (0))\cdot(1+\cos (0))}{\cos (0)}=2"


The tangent line in point-slope form


"y-0=2(x-0)"

The tangent line in slope-intercept form


"y=2x"



"t_2=\pi:"


"slope_2=\dfrac{\cos((\pi)+\sin (0))\cdot(1+\cos (\pi))}{\cos (\pi)}=0"

The tangent line in point-slope form


"y-0=0(x-0)"

The tangent line in slope-intercept form


"y=0"



2.


"\begin{matrix}
 x=-1 \\
 y=-1
\end{matrix}, \ 0\leq t\leq \pi"

"\begin{matrix}
 2 \cos t=-1 \\
 2-\cos^2t=-1
\end{matrix}, \ \ 0\leq t\leq \pi"

No solution. Point "(-1, -1)" does not lie on the given curve.

"slope=\dfrac{dy}{dx}=\dfrac{dy/dt}{dx/dt}=\dfrac{2\cos t\sin t}{-2\sin t}=-\cos t,"

"0<t<\pi"

Point "(1, 1.75)"


"\begin{matrix}
 x=1 \\
 y=1.75
\end{matrix}, \ 0\leq t\leq \pi"

"\begin{matrix}
 2\cos t=1 \\
 2-\cos^2 t=1.75
\end{matrix}, \ 0\leq t\leq \pi"

"t=\pi/3""slope=\cos(\pi/3)=-0.5"


The tangent line in point-slope form


"y-1.75=-0.5(x-1)"

The tangent line in slope-intercept form


"y=-0.5x+2.25"